# A Scalable Optimization Approach for Equitable Facility Location: Methodology and Transportation Applications

### v3

@@ 3.4.1 Mathematical Formulation
It is rarely the case that all potential locations are equally suitable. For example, in the equitable polling locations application, early voting sites are open on many days leading up to an election and this can be disruptive to normal operations at the selected sites. When this model was used by advocacy groups in preparation for the 2024 United States presidential election, see code base of Skipper et al. (2025), fire stations and churches were supplied as potential early voting sites in one Georgia county, but were only to be selected if their inclusion would be particularly impactful with respect to voter access. Similarly, if an urban planning board for a particular city were selecting supermarket locations to improve food access, some potential locations would be less suitable than others due to land costs or other practical considerations.
Suppose $U\subseteq S$ is a set of less-desirable potential locations. For each $s\in U$, let $c_{s}\geq 0$ represent the number of units of improvement in the Kolm-Pollak score that would need to be realized (versus an optimal score not using site $s$) to select location $s$. We want to incorporate these preferences in the (KPL) model by penalizing the optimal Kolm-Pollak score by $\sigma:=\sum_{s\in U}c_{s}x_{s}$ units. However, we cannot do this directly because the objective function, (8), does not minimize the Kolm-Pollak score directly. The following result provides the value that the (KPL) linear objective function, $\overline{\mathcal{K}}$, must be penalized by for the associated Kolm-Pollak score, $\mathcal{K}$, to be penalized by $\sigma$ meters.
@@
[…]
@@
The accuracy of the penalty term depends both on the quality of the approximation $\hat{\mathcal{K}}$ in the objective function (12) and on the linearization points, $\boldsymbol{\beta}$, chosen for (15). In the rest of this section, we analyze the error to the penalty term introduced by each of these approximations, which leads to guidance on choosing $\hat{\mathcal{K}}$ and $\boldsymbol{\beta}$ to keep this error low.
@@ 3.4.2 Error Analysis and Implementation Guidelines
1. Penalty error introduced by parameter approximation:
Let $\sigma^{*}:=\sum_{s\in U}c_{s}x_{s}^{*}\geq 0$ represent the penalty associated with stores selected by optimal solution $(\mathbf{x}^{*},\mathbf{y}^{*})$ to (KPLp). Let $\hat{\sigma}$ represent the actual penalty (approximating $\sigma^{*}$) that is applied to an optimal Kolm-Pollak score by (KPLp). Let $\Delta:=\hat{\mathcal{K}}-\mathcal{K}^{*}$ represent the signed error in approximating $\mathcal{K}^{*}$ by $\hat{\mathcal{K}}$.
@@ 1. Penalty error introduced by parameter approximation:
Proof.
$=~{}-\frac{1}{\kappa}\ln\left(e^{-\kappa\mathcal{K}^{*}}+e^{-\kappa\hat{\mathcal{K}}-\kappa\sigma^{*}}-e^{-\kappa\hat{\mathcal{K}}}\right)$
so that,
@@
Proof.
(1) Suppose $\sigma^{*}>0$ is fixed. We have
Recalling that $\kappa<0$, we have that $e^{-\kappa\sigma^{*}}>1$, so that $\frac{\delta}{\delta\Delta}\hat{E}(\Delta,\sigma^{*})\in(0,1)$. Thus, $\hat{E}(\Delta,\sigma^{*})$ is increasing (though not quickly) in $\Delta$. It is straightforward to evaluate $\hat{E}(0,\sigma^{*})=0$. Thus, when $\Delta>0$, $\hat{E}(\Delta,\sigma^{*})>0$ and $\hat{E}(\Delta,\sigma^{*})=|\hat{E}(\Delta,\sigma^{*})|$ is increasing in $|\Delta|$. Similarly, when $\Delta<0$, $\hat{E}(0,\sigma^{*})<0$ and $-\hat{E}(\Delta,\sigma^{*})=|\hat{E}(\Delta,\sigma^{*})|$ is increasing in $|\Delta|$.
@@
Proof.
Thus,
Finally,
$~{}=~{}\mathcal{K}^{*}+\sigma^{*}+\Delta$
$~{}\leq~{}\mathcal{K}^{*}+\sigma^{*}+\hat{E}(\Delta,\sigma^{*})$
$~{}\leq~{}\mathcal{K}^{all}+\sigma^{all},$
which means that $\sigma^{*}\leq\sigma^{all}$, as required. ∎
2. Penalty error introduced by linearization:
Next we analyse the error in the penalty term introduced by approximating $e^{q}$, $q=-\kappa\sum_{s\in U}c_{s}x_{s}=-\kappa\sigma$, via lower-bounding tangent lines in (KPLt). Let $\ddot{\sigma}$ represent the penalty applied to an optimal Kolm-Pollak score by (KPLt). Thus, $\ddot{\sigma}$ is an under-approximation of $\hat{\sigma}$ from the previous subsection.
@@ 2. Penalty error introduced by linearization:
$g(q):=\begin{cases}1+q,&\text{for }q\in[0,q_{0,1});\\ e^{\beta_{i}}+e^{\beta_{i}}(q-\beta_{i}),&\text{for }q\in[q_{i-1,i},q_{i,i+1}),i=1,2,3,\dots,n-1;\\ e^{\beta_{n}}+e^{\beta_{n}}(q+\beta_{n})&\text{for }q\in[q_{n-1,n},-\kappa\sigma^{max}],\\ \end{cases}$
@@
In practice, the penalties, $c_{s}$, would often be the same, or would only take on a few different values. If the penalties are all the same ($c_{s}=c,\forall s\in U$), we can choose linearization points so that (KPLt) is an exact formulation of (KPLp).
@@
Proof.
Due to (1e), for all feasible $\mathbf{x}:=(x_{1},x_{2},\dots,x_{|S|})$, $q=-\kappa\sum_{s\in U}cx_{s}$ can only take the values of the selected linearization points. At these values of $q$, the approximating tangent lines match the exponential expression, $e^{q}$, exactly. ∎
When it is not efficient to choose a linearization point for every possible value of $q~{}=~{}-\kappa\sum_{s\in U}c_{s}x_{s}$ (i.e., when there are more than a few unique penalty values), we would like to choose linearization points $\beta_{i}$ so that the approximation error is small. Tangent lines do not closely approximate exponential functions in general (Proposition 12 and Corollary 13). However, the linearization provides a very close approximation in this application (Proposition 14 and Theorem 16).
@@
Thus, if $a$ is large, choosing $w$ small enough to have a reasonably small error between $e^{x}$ and its lower-bounding tangent lines will introduce numerical problems in a computational setting. However, in practical instances of (KPLq), the argument of the exponential expression in the penalty expression, $q=-\kappa\sum_{s\in U}c_{s}x_{s}$, is bounded above by $|\epsilon|$, so that we can achieve a very close approximation to $e^{q}$ with a numerically stable choice of $w$. (Recall that $\epsilon$, the aversion to inequality parameter, is negative and typically takes values between $-\frac{1}{2}$ and $-2$.)
@@
Proof.
where the final inequality is due to the Cauchy-Schwarz inequality. ∎
@@
Proposition 14 indicates that the exponential expression $e^{q}$ can be closely approximated by tangent lines in (KPLt) and we can derive bounds directly on the error in the Kolm-Pollak penalty that is introduced by the tangent line approximation. We assume that the linearization points are equally spaced at a width of $w<1$ apart. For context with regards to (18), $A(0.01)\approx 10^{-5}$ and $A(0.001)\approx 10^{-7}$.
@@
Proof.
$~{}\geq~{}0.$
We have $\ddot{\sigma}\leq\hat{\sigma}$ because $err(q^{*},w)\geq 0$, and for (17),
For the first part or (18), if $\Delta\leq 0$, then $1+e^{\kappa\sigma^{*}}(e^{\kappa\Delta}-1)\geq 1$, so that $\hat{\sigma}-\ddot{\sigma}\leq\frac{1}{\kappa}\ln\left(1-A(w)\right)$. For the second part of (18), if $\hat{\mathcal{K}}=\mathcal{K}^{rem}$, then $\mathcal{K}^{*}+\sigma^{*}\leq\mathcal{K}^{*}+\hat{\sigma}\leq\mathcal{K}^{rem}+0$, where the first inequality is due to Theorem 5 and the second is due to the optimality of (KPLp). Thus, $0\leq\sigma^{*}\leq\mathcal{K}^{rem}-\mathcal{K}^{*}=\Delta$, and we have,
$0.751+e^{\kappa\sigma^{*}}(e^{\kappa\Delta}-1)~{}\geq~{}1+e^{\kappa\Delta}(e^{\kappa\Delta}-1)~{}\geq~{}$ 0.75 .
The following is a direct result of Theorem (10), Proposition (11), and Theorem (16).
@@
(Penalty error bounds) We solved four realistic instances of (KPLt) using the data for Santa Rosa, California from the food desert application described in detail in Section 4. In all instances, the linearization points were chosen to be a fixed width apart, differing only in the width, $w$. We set $k=10$, $\epsilon=-1$, $\kappa=-\alpha=-6.93\times 10^{-4}$, and $\hat{\mathcal{K}}=\mathcal{K}^{all}\approx 1273.5382$ m (meters). We chose $U$ to be the set of locations optimal to (KPL)all and $c_{s}=c=4.8609\approx\frac{\mathcal{K}^{rem}-\mathcal{K}^{all}}{10}$, for all $s\in U$, so that $\sigma^{all}=48.609$ m. The optimal locations selected by (KPLt) were the same for all instances and included two penalized locations, as displayed in Figure 2. For all instances, $\mathcal{K}^{*}\approx 1280.1363$ m, $\Delta=\hat{\mathcal{K}}-\mathcal{K}^{*}\approx-6.5981$ m, and $\sigma^{*}=9.7217$ m. Note that $-\kappa\sigma^{max}\approx 0.0337$, so we explore values of $w$ as large as $0.01$.
Tables 5 and 6 contain bounds on the error in the Kolm-Pollak penalty introduced by approximating $\hat{\mathcal{K}}$ and by the linear approximations, respectively, in (KPLt). Table 7 contains combined error bounds, as well as the actual observed separate and combined errors for each instance.
@@ 3.4.3 Penalty Strategy and Example
We conclude this section by summarizing strategies for modeling the penalty term recommended by our analysis, then demonstrating their practical implementation. Choose $\hat{\mathcal{K}}=\mathcal{K}^{all}$ to keep the error associated with approximating $\hat{\mathcal{K}}$ low. A good way to scale the per-location penalty so that penalized locations are discouraged but not eliminated is to set
@@
where $N$ represents the number of less-desirable locations selected by (KPL)all. If all per-location penalties are the same ($c$), set $w:=-\kappa c$ to avoid any linearization error. Otherwise, scale $w$ according to the $\Delta\leq 0$ case in (18) to bring the linearization error below a desired threshold.

### v4

@@ 3.3 Incorporating Location Preferences Through Penalties
Note that for clarity of presentation, proofs of all theorems and propositions given in this section can be found in A.
It is rarely the case that all potential locations are equally suitable. For example, in the equitable polling locations application, early voting sites are open on many days leading up to an election and this can be disruptive to normal operations at the selected sites. When this model was used by advocacy groups in preparation for the 2024 United States presidential election, fire stations and churches were supplied as potential early voting sites in one Georgia county, but were only to be selected if their inclusion would be particularly impactful with respect to voter access (Skipper et al., 2025). Similarly, if an urban planning board for a particular city were selecting supermarket locations to improve food access, some potential locations would be less suitable than others due to land costs or other practical considerations.
@@ 3.3.1 Mathematical Formulation
Suppose $U\subseteq S$ is a set of less-desirable potential locations. For each $s\in U$, let $c_{s}\geq 0$ represent the number of units of improvement in the Kolm-Pollak score that would need to be realized (versus an optimal score not using site $s$) to select location $s$. We want to incorporate these preferences in the (KPL) model by penalizing the optimal Kolm-Pollak score by $\sigma:=\sum_{s\in U}c_{s}x_{s}$ units. However, we cannot do this directly because the objective function, given by (8), does not minimize the Kolm-Pollak score directly. The following result provides the value that the (KPL) linear objective function, $\overline{\mathcal{K}}$, must be penalized by for the associated Kolm-Pollak score, $\mathcal{K}$, to be penalized by $\sigma$ meters.
@@
In the objective function (13), $\hat{\mathcal{K}}$ is a parameter approximating $\mathcal{K}^{*}$, the unpenalized Kolm-Pollak score associated with an optimal solution to (KPLp): $\mathcal{K}^{*}:=-\frac{1}{\kappa}\ln\left(\frac{1}{T}\overline{\mathcal{K}}(\mathbf{y}^{*})\right)$, where $(\mathbf{x}^{*},\mathbf{y}^{*})$ is optimal to (KPLp). Moreover, to use a linear optimization solver, we must linearize the one-dimensional exponential function in constraint (14). Let $\sigma^{max}$ represent the largest possible value that $\sigma=\sum_{s\in U}c_{s}x_{s}$ can take for a feasible solution, $(\mathbf{x},\mathbf{y})$, to (KPLp). We can accomplish the linearization by replacing constraint (14) with a set of lower-bounding tangent lines constructed at points $(q,e^{q})$, for $q\in\boldsymbol{\beta}=\{\beta_{0},\beta_{1},\dots,\beta_{n}\}$, where $0=\beta_{0}<\beta_{1}<\dots<\beta_{n-1}<-\kappa\sigma^{max}\leq\beta_{n}$, as follows:
@@
The accuracy of the penalty term depends both on the quality of the approximation $\hat{\mathcal{K}}$ in the objective function (13) and on the linearization points, $\boldsymbol{\beta}$, chosen for constraint (16). In the rest of this section, we analyze the error to the penalty term introduced by each of these approximations, which leads to guidance on choosing $\hat{\mathcal{K}}$ and $\boldsymbol{\beta}$ to keep this error low.
@@ 3.3.2 Error Analysis and Implementation Guidelines
First we consider the penalty error introduced by parameter approximation. Let $\sigma^{*}:=\sum_{s\in U}c_{s}x_{s}^{*}\geq 0$ represent the penalty associated with stores selected by optimal solution $(\mathbf{x}^{*},\mathbf{y}^{*})$ to (KPLp). Let $\hat{\sigma}$ represent the actual penalty (approximating $\sigma^{*}$) that is applied to an optimal Kolm-Pollak score by (KPLp). Let $\Delta:=\hat{\mathcal{K}}-\mathcal{K}^{*}$ represent the signed error in approximating $\mathcal{K}^{*}$ by $\hat{\mathcal{K}}$.
@@
Next we analyze the penalty error introduced by linearization, i.e., by approximating $e^{q}$, $q=-\kappa\sum_{s\in U}c_{s}x_{s}=-\kappa\sigma$, via lower-bounding tangent lines in (KPLt). Let $\ddot{\sigma}$ represent the penalty applied to an optimal Kolm-Pollak score by (KPLt). Thus, $\ddot{\sigma}$ is an under-approximation of $\hat{\sigma}$.
@@
[…]
@@
In practice, the penalties, $c_{s}$, would often be the same, or would only take on a few different values. If the penalties are all the same (i.e., $c_{s}=c,\forall s\in U$), we can choose linearization points so that (KPLt) is an exact formulation of (KPLp).
@@
When it is not efficient to choose a linearization point for every possible value of $q~=~-\kappa\sum_{s\in U}c_{s}x_{s}$ (i.e., when there are more than a few unique penalty values), we would like to choose linearization points $\beta_{i}$ so that the approximation error is small. Tangent lines do not closely approximate exponential functions in general (Proposition 12 and Corollary 13). However, the linearization does provide a very close approximation in this application (Proposition 14 and Theorem 16).
@@
Thus, if $a$ is large, choosing $w$ small enough to have a reasonably small error between $e^{x}$ and its lower-bounding tangent lines will introduce numerical problems in a computational setting. However, in practical instances of (KPLq), the argument of the exponential expression in the penalty expression, $q=-\kappa\sum_{s\in U}c_{s}x_{s}$, is bounded above by $|\epsilon|$, so that we can achieve a very close approximation to $e^{q}$ with a numerically stable choice of $w$. (Recall that $\epsilon$, the aversion to inequality parameter, is negative and typically takes values between $\frac{1}{2}$ and 2-2.)
@@
Proposition 14 indicates that the exponential expression $e^{q}$ can be closely approximated by tangent lines in (KPLt) and we can derive bounds directly on the error in the Kolm-Pollak penalty that is introduced by the tangent line approximation. We assume that the linearization points are equally spaced at a width of $w<1$ apart. For context with regards to inequality (19), $A(0.01)\approx 10^{-5}$ and $A(0.001)\approx 10^{-7}$.
@@
The following is a direct result of Theorem 10, Proposition 11, and Theorem 16.
@@
(Penalty error bounds) To explore the error observed in practice, we solved four realistic instances of (KPLt) using the data for Santa Rosa, California from the food desert application described in Section 5.1.1. In all instances, the linearization points were chosen to be a fixed width apart, $w$. We set $k=10$, $\epsilon=-1$, $\kappa=-\alpha=-6.93\times 10^{-4}$, and $\hat{\mathcal{K}}=\mathcal{K}^{all}\approx 1273.5382$ m (meters). We chose $U$ to be the set of locations optimal to (KPL)all and $c_{s}=c=4.8609\approx\frac{\mathcal{K}^{rem}-\mathcal{K}^{all}}{10}$, for all $s\in U$, so that $\sigma^{all}=48.609$ m. The optimal locations selected by (KPLt) were the same for all instances and included two penalized locations, as displayed in Figure 3. For all instances, $\mathcal{K}^{*}\approx 1280.1363$ m, $\Delta=\hat{\mathcal{K}}-\mathcal{K}^{*}\approx-6.5981$ m, and $\sigma^{*}=9.7217$ m. Note that $\kappa\sigma^{max}\approx 0.0337$, so we explore values of $w$ as large as $0.01$.
Tables 3 and 4 contain bounds on the error in the Kolm-Pollak penalty introduced by approximating $\hat{\mathcal{K}}$ and by the linear approximations, respectively, in (KPLt). Table 5 contains combined error bounds, as well as the actual observed separate and combined errors for each instance.
@@ 3.3.3 Penalty Strategy and Example
We conclude this section by summarizing strategies for modeling the penalty term recommended by our analysis, then demonstrating their practical implementation.
@@
where $N$ represents the number of less-desirable locations selected by (KPL)all. If all per-location penalties are the same ($c$), set $w:=-\kappa c$ to avoid any linearization error. Otherwise, scale $w$ according to the $\Delta\leq 0$ case in inequality (19) to bring the linearization error below a desired threshold.
@@
Table 6: The Kolm-Pollak scores ($\mathcal{K}$) and the number of fire stations and churches ($|P|$) selected by optimal solutions to three models in the Gwinnett County early voting application. Note that (KPLrem) chooses no less-suitable sites, but leads to the highest Kolm-Pollak score. (KPLall) achieves the best possible Kolm-Pollak scores, while (KPLtt) achieves near-optimal Kolm-Pollak scores with fewer less-suitable sites, as measured by “gap”: $\frac{\mathcal{K}^{t}-\mathcal{K}^{all}}{\mathcal{K}^{all}}$. The Kolm-Pollak scores, the per-site desired penalty, and the per-site penalty error are all measured in meters.
4 Scaling the Inequality Aversion Parameter
To minimize the Kolm-Pollak EDE via (KPL), we must approximate $\alpha$, the factor used to scale the inequality aversion parameter. As a consequence, we do not have complete control over the aversion to inequality represented by the optimal Kolm-Pollak score, however, we are able to precisely quantify the effect this has. In this section, we explore computationally the impact of approximating $\alpha$ on both the aversion to inequality represented by the optimal distribution and on the optimal Kolm-Pollak score itself. Our tests include practical instances on 428 cities in which we seek to optimally locate 1, 5, or 10 additional supermarkets, as well as 5 cities in which we replace all existing polling sites with the same number of optimally located sites.
4.1 Iterative Calibration of $\alpha$
Let $\epsilon_{0}$ represent the desired level of aversion to inequality. For each problem instance, we first estimate the optimal distribution of distances by assigning each census block to the closest existing location (grocery store or polling site). We use this distribution of distances to calculate an approximate value of $\alpha$, which we denote $\alpha_{1}^{in}$. Let (KPL)1 represent the model (KPL) with $\kappa=\kappa_{1}:=\alpha_{1}^{in}\epsilon_{0}$. Let $\alpha_{1}^{out}$ denote the value of $\alpha$ associated with the optimal solution to (KPL)1. We can calculate the aversion to inequality represented by the optimal solution to (KPL)1 as $\epsilon_{1}:=\kappa_{1}/\alpha_{1}^{out}$. Comparing $\epsilon_{0}$ and $\epsilon_{1}$ is a useful way to check the accuracy of $\alpha_{1}^{in}$. The closer $\alpha_{1}^{in}$ approximates $\alpha$, the closer $\epsilon_{1}$ will be to $\epsilon_{0}$.
It is reasonable to expect that the distribution of distances associated with an optimal solution to (KPL)1 will produce a more accurate $\alpha$-estimate. Thus, we solve (KPL) again, this time with $\kappa=\kappa_{2}:=\epsilon_{0}\alpha_{2}^{in}$, where $\alpha_{2}^{in}:=\alpha_{1}^{out}$. We denote this version of (KPL) as (KPL)2. We compute the aversion to inequality associated with the optimal solution to (KPL)2 as $\epsilon_{2}:=\kappa_{2}/\alpha_{2}^{out}$, where $\alpha_{2}^{out}$ is calculated using the optimal distribution of distances obtained by solving (KPL)2.
In extensive computational tests, we found that the improved approximation, $\alpha_{2}^{in}$, was extremely accurate (results provided below). However, we also want to understand how important it is tune the $\alpha$ parameter in this way. In other words, we want to test how sensitive optimal solutions are to the quality of the $\alpha$ approximation to determine if it is worth the extra computation time to solve (KPL) twice to obtain and use the very accurate $\alpha_{2}^{in}$. To do this, we used optimal solutions to (KPL)2 as a benchmark to test the quality of optimal solutions to (KPL)1. We denote the Kolm-Pollak EDE associated with an optimal solution to (KPL)n as $\mathcal{K}_{n}^{*}$, for $n\in\{1,2\}$. $\mathcal{K}_{n}^{*}$ represents the “true” value of the metric, calculated exactly using $\kappa=\epsilon_{0}\alpha_{n}^{out}$. We find that $\mathcal{K}_{1}^{*}$ very closely approximates the more accurate $\mathcal{K}_{2}^{*}$ (detailed results below).
4.2 $\alpha$-Approximation Experiments: Grocery Stores
As described in detail in Section 5, the bulk of our computational experiments involved optimally locating 1, 5, and 10 additional grocery stores in the 500 largest cities in the United States. We completed two sets of experiments, one using a high-performance computing cluster and another with a personal laptop (specifications given in Section 5.1.2). For the $\alpha$-approximation analysis described here, we used the laptop, which did not have sufficient memory to load the data for the largest cities. Thus, this study includes the 428 cities for which the laptop was able to solve (KPL)1 and (KPL)2 for 1, 5, and 10 additional stores.
Figure 5 presents the distributions of $\epsilon_{1}$ (top row) and $\epsilon_{2}$ (bottom row) for the 428 cities when locating 1, 5, and 10 new stores. We see that $\epsilon_{1}$ has a skewed Normal distribution with a mean value that represents less aversion to inequality than $\epsilon_{0}=-1$ (which is indicated by the red vertical line). As the number of new stores increases, the center of the $\epsilon_{1}$ distribution moves farther from the desired value of 1-1. This shift makes sense intuitively because adding more new stores naturally results in greater changes to the distribution of distances as compared to the original distribution, which was used in the original $\alpha$ approximation, $\alpha_{1}^{in}$.
Figure 5: Distribution of the “true” value of $\epsilon$, after the initial run ($\epsilon_{1}$) and after the second run ($\epsilon_{2}$), for $k=1,5,\text{ and }10$ new stores placed in 428 US cities. The desired value is $\epsilon=-1$ and is indicated by the red vertical line.
As expected, $\alpha_{1}^{out}$ is a very good approximation, and $\epsilon_{2}$ very closely approximates $\epsilon_{0}$ in every instance. We can conclude that (KPL)2 achieves a high quality optimal solution with respect to the desired level of inequality aversion. In order to test the quality of the solutions achieved by (KPL)1, we calculated the raw gap between $\mathcal{K}_{1}^{*}$ and $\mathcal{K}_{2}^{*}$ as $\Delta:=|\mathcal{K}_{1}^{*}-\mathcal{K}_{2}^{*}|$ meters, and the percent gap as gap $\Delta/\mathcal{K}_{2}^{*}$. These results are visualized in Figure 6, where the chart displays the cumulative distribution of gap for the 428 cites, separated out by the number of additional grocery stores.